6) Suppose the working-age population is 220 million, the labor force is 150 million, and the unemployment rate is 10 percent. The number of unemployed people is

A) 7 million

B) 15 million.

C) 22 million

D) 37 million.

E) 70 million.
Answer #1

Answer: B) 15 million.

Unemployment rate = (Total number of people unemployed / Total number of people in labor force) * 100

Total number of people unemployed =(Unemployment rate * Total number of people in labor force)/100

Here,  labor force = 150 million ; unemployment rate = 10%

Total number of people unemployed = (10 * 150) million /100 = 15 million.

So, the number of unemployed people is 15 million.

[ 150 million = 150,000,000

Total number of people unemployed =(150,000,000 *10) million/ 100 = 15,000,000 = 15 million]

___________________________________________________